OpenStudy (anonymous):

Find the derivative of y with respect to x,t or theta as appropriate. (a) y= ln(x)/x^6 (b) y= ln((1-x)/(x+5)^3)

OpenStudy (aravindg):

(a) use quotient rule

OpenStudy (anonymous):

what do you do after that

OpenStudy (aravindg):

simplify if needed

OpenStudy (anonymous):

thanks

OpenStudy (aravindg):

yw :)

OpenStudy (anonymous):

i got (x^6(ln) - ln(x) (6x^5))/x^12 the answer is 1-6ln(x)/x^7 im stuck with simplifying

OpenStudy (anonymous):

i would always change those equations into product rule

OpenStudy (anonymous):

\[y = \ln(x).x^{-6}\]

OpenStudy (anonymous):

\[=x^{-1}x^{-6}-6x^{-7}\ln(x)\]

OpenStudy (anonymous):

\[=x^{-7}-6x^{-7}\ln(x)=x^{-7}(1-6\ln(x))\]

OpenStudy (anonymous):

as for the 2nd problem the answer would be

OpenStudy (anonymous):

=[-1/(1-x)(x+5)^3] -[3ln(1-x)/(x+5)^4]

OpenStudy (anonymous):

thank you